OpenStudy (anonymous):

Marks: 6 [Part 1 (4 points)] Show all work in simplifying the quantity 14 x squared minus 123 x plus 34 all over 2 x squared minus 13 x minus 34 (list restrictions). [Part 2 (2 points)] Use complete sentences to explain your process.

OpenStudy (anonymous):

\[14x^2-123x+34/2x^2-13x-34\]

OpenStudy (campbell_st):

\[\frac{14x^2 - 123x + 34}{2x^2 - 13x -34}\] factorise the denominator for the restrictions (2x-17)(x+2) = 0 solve 2x - 17 = 0 and x + 2 = 0 to find the restrictions

OpenStudy (anonymous):

So thats all you do for the restrictions? How do you simplify the problem

OpenStudy (campbell_st):

the numerator can be factorised as well so you have \[\frac{(2x -17)(7x - 2)}{(2x-17)(x+2)}\] as (2x-17) is a common factor the solution to 2x-17 is a point of discontinuity the simplification should be obvious

OpenStudy (anonymous):

so the answer would be 7x-2/x+2